# FaithUn: Toward Faithful Forgetting in Language Models by Investigating the Interconnectedness of Knowledge

### v1

@@ Abstract
Various studies have attempted to remove sensitive or private knowledge from a language model to prevent its unauthorized exposure. However, prior studies have overlooked the complex and interconnected nature of knowledge, where related knowledge must be carefully examined. Specifically, they have failed to evaluate whether an unlearning method faithfully erases interconnected knowledge that should be removed, retaining knowledge that appears relevant but exists in a completely different context. To resolve this problem, we first define a new concept called superficial unlearning, which refers to the phenomenon where an unlearning method either fails to erase the interconnected knowledge it should remove or unintentionally erases irrelevant knowledge. Based on the definition, we introduce a new benchmark, FaithUn, to analyze and evaluate the faithfulness of unlearning in real-world knowledge QA settings. Furthermore, we propose a novel unlearning method, KLUE, which updates only knowledge-related neurons to achieve faithful unlearning. KLUE identifies knowledge neurons using an explainability method and updates only those neurons using selected unforgotten samples. Experimental results demonstrate that widely-used unlearning methods fail to ensure faithful unlearning, while our method shows significant effectiveness in real-world QA unlearning.
@@
Nakyeong Yang1, Minsung Kim1, Seunghyun Yoon2, Joongbo Shin3 and Kyomin Jung1 1Seoul National University, 2Adobe Research, 3LG AI Research {yny0506, kms0805, kjung}@snu.ac.kr syoon@adobe.com, jb.shin@lgresearch.ai
@@ 1 Introduction
Large language models (LLMs) are trained on a vast corpus of text, enabling them to achieve outstanding performance across various tasks (Radford et al., 2019; Chowdhery et al., 2023; Gemma et al., 2024). However, LLMs may present privacy risks, as sensitive or private information could unintentionally be included in the large text corpus used for training. Therefore, prior studies have investigated unlearning undesirable knowledge in language models. To assess unlearning, most studies have examined whether a model successfully forgets the targeted knowledge while retaining irrelevant knowledge (Shi et al., 2024; Li et al., 2024a; Maini et al., 2024; Jin et al., 2024).
@@
Figure 1: Faithful Unlearning. FaithUn proposes three types of datasets to evaluate the faithfulness of unlearning methods (i.e., Paraphrased, Multi-hop, and Same-answer datasets). Each target knowledge to be unlearned is mapped with questions corresponding to these three dataset types for evaluation.
However, they are limited since they have overlooked the complex and interconnected nature of knowledge, which requires careful investigation of related knowledge. Specifically, these studies have examined only the independent knowledge and failed to evaluate whether an unlearning method effectively erases interconnected knowledge that should be removed, while retaining knowledge that appears relevant but exists in a completely different context. Figure 1 presents an example of faithful unlearning in the real-world knowledge setting. Unlearning methods should also remove paraphrased and multi-hop questions, as they involve knowledge interconnected with the target question being unlearned. Conversely, unlearning methods should retain knowledge of other questions with the same answer as the target, if they actually contain different knowledge despite appearing relevant.
To address this gap, we first define superficial unlearning, which refers to the phenomenon where an unlearning method either fails to erase the interconnected knowledge it should remove or unintentionally erases irrelevant knowledge. Based on the definition, we introduce FaithUn (A Faithful Unlearning Evaluation Benchmark for Real-world Knowledge Question Answering), a new benchmark to examine three types of unlearning challenges: generalization, multi-hop unlearning, and shortcut unlearning to investigate superficial unlearning. Generalization (Anil et al., 2022; Yang et al., 2024a; Albalak et al., 2024), the multi-hop reasoning (Zhong et al., 2023; Li et al., 2024b; Yang et al., 2024b), and shortcut learning (Du et al., 2023; Tang et al., 2023; Zhou et al., 2023) are crucial challenges in machine learning research. Since the unlearning process typically relies on fewer data instances than general training, these challenges can be further amplified. Therefore, we construct three types of new datasets—paraphrased, multi-hop, and same-answer datasets—to examine superficial unlearning. These datasets address generalization, multi-hop knowledge unlearning, and shortcut unlearning, respectively. We demonstrate that existing unlearning methods do not ensure faithful unlearning, which raises new research questions for knowledge unlearning.
Furthermore, we propose a novel method, KLUE, which stands for Knowledge-Localized UnlEarning, to achieve faithful unlearning by precisely identifying and updating neurons related to the target knowledge. Specifically, we use attribution (Yang et al., 2023), an explainability method, to determine which neurons should be updated by quantifying how much information each neuron contributes to predicting the answer to a given question. However, the quantified score may include superficial knowledge that simply affects the target output’s probability without considering contextual meaning. Therefore, we propose a robust knowledge regularization method that accurately quantifies each neuron’s knowledge score, mitigating the superficial contribution of neurons. After identifying knowledge neurons, our method selectively unlearns the target knowledge while preserving other knowledge by updating only knowledge-related neurons with selected unforgotten samples. In our experiments, our method significantly outperforms the baselines in the FaithUn setting, demonstrating that knowledge-localized unlearning effectively achieves faithful unlearning.
2 Unlearning in Large Language Models
Machine unlearning has been used as a solution to address privacy and copyright issues in the text generation process of language models. Notable examples include gradient ascent-based methods (Jang et al., 2023; Yao et al., 2023; Barbulescu and Triantafillou, 2024), preference optimization approaches (Rafailov et al., 2024; Zhang et al., 2024; Jin et al., 2024), and representation learning techniques (Li et al., 2024a; Yao et al., 2024).
However, the effectiveness of these methods has not been clearly demonstrated, prompting prior studies to introduce benchmarks in the field of unlearning to assess them. Eldan and Russinovich (2023); Shi et al. (2024); Tian et al. (2024) have aimed to unlearn the knowledge of copyrighted texts (e.g., BBC News and Harry Potter book) in a language model. Li et al. (2024a) have introduced a benchmark dealing with hazardous knowledge in various professional domains (e.g., biosecurity and cybersecurity). Maini et al. (2024); Jin et al. (2024) have proposed benchmarks for unlearning various entities. Specifically, Maini et al. (2024) have created synthetic entity profiles and removed their knowledge from a language model. Jin et al. (2024) have tried to unlearn the knowledge about real-world entities and evaluated the knowledge memorization in various forms of assessment (e.g., cloze test and question answering). However, existing studies remain limited as they have only examined independent knowledge and overlooked the intricate nature of world knowledge. World knowledge is highly complex and interconnected, which means that unlearning the target knowledge requires examining related knowledge carefully. Our research focuses on this aspect, examining and facilitating faithful unlearning.
3 The FaithUn Benchmark
3.1 Problem Definition
The FaithUn task evaluates unlearning algorithms under real-world knowledge QA settings. Formally, given a language model $P_{\theta}(y|x)=\prod_{t=1}^{T}P_{\theta}(y_{t}|x,y_{1},...,y_{t-1})$ with parameters $\theta$, an unlearning algorithm $f$ updates $\theta$ to $\theta^{\prime}$, erasing the target knowledge from $P_{\theta}$. FaithUn includes various question-answer pairs $(q,a)\in\mathcal{C}$, where $\mathcal{C}$ is a question-answer pair set. Our task provides forget set $\mathcal{C}_{f}$, which contains target question-answer pairs to be forgotten, where $\mathcal{C}_{f}\subset\mathcal{C}$. FaithUn also provides retain set $\mathcal{C}_{r}\subset\mathcal{C}\backslash\mathcal{C}_{f}$ and test set $\mathcal{C}_{t}\subset\mathcal{C}\backslash(\mathcal{C}_{f}\cup\mathcal{C}_{r})$. $\mathcal{C}_{r}$ is used in the unlearning process as training samples to maintain the original knowledge of $P_{\theta}$, and $\mathcal{C}_{t}$ is used as unseen data to evaluate an unlearned model $P_{\theta^{\prime}}$ to reveal whether the unlearned model maintains the original knowledge. Furthermore, FaithUn provides other new types of test sets (i.e., paraphrased, multi-hop, and same-answer sets) to assess the faithfulness of unlearning methods. Before introducing the other datasets, we first define key aspects of our benchmark.
@@ 3.1 Problem Definition
Knowledge Exists in LLMs X X O X O O
Generalization Test X X X X O O
Multi-hop Unlearning Test X X X X X O
Shortcut Unlearning Test X X X X X O
@@
Table 1: Dataset Comparison. FaithUn aims to examine three types of unlearning challenges: generalization, multi-hop knowledge unlearning, and shortcut unlearning to investigate superficial unlearning. FaithUn can be used flexibly as it evaluates the removal of pre-existing knowledge about famous figures within LLMs.
@@
A world knowledge graph $\mathcal{K}$ is a directed multi-graph where nodes are entities and edges are labeled with relations, i.e., elements of two sets $\mathcal{E}$ and $\mathcal{R}$, respectively. We define $\mathcal{K}$ as a collection of triples $(s,r,o)\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$, where $s$, $r$, $o$ denote the subject, relation, and object, respectively (Ruffinelli et al., 2020; Loconte et al., 2024). We assume that a world knowledge question is mapped to triples of $\mathcal{K}$; thus, we also define a knowledge mapping function, $\tau:\mathcal{Q}\rightarrow\mathcal{P}(\mathcal{K})$, where $\mathcal{Q}$ is a set of questions and $\mathcal{P}(\mathcal{K})$ represents the power set of $\mathcal{K}$. For example, the knowledge of a multi-hop question, $q_{i}$ = "Which continent is Tom Cruise’s country in?", can be denoted as a set of triples like $\kappa_{i}=\tau(q_{i})$ = {("Tom Cruise", "country", "U.S.A"), ("U.S.A", "continent", "North America")}.
@@
where $\iota$ is an input prompt template for the language model $P_{\theta}$, and $\mathcal{Q}$ and $\mathcal{A}$ are question and answer sets, respectively. From the definition, $\mathcal{M}_{\theta}(q,a)=1$ indicates that the language model retains the knowledge of $(q,a)$, while $\mathcal{M}_{\theta}(q,a)=0$ signifies that it does not.
@@
where $\kappa_{f}=\bigcup_{(q,a)\in\mathcal{C}_{f}}{\tau(q)}$ and $\kappa_{j}=\tau(q_{j})$.
For example, suppose that an unlearning algorithm $g$ unlearns the knowledge of the question $q_{i}$ = "Which country is Tom Cruise from?", but it does not unlearn the multi-hop question $q_{j}$ = "Which continent is Tom Cruise’s country in?". Then, the knowledge of two questions can be denoted as a set of knowledge triples like $\kappa_{i}=\tau(q_{i})=$("Tom Cruise", "country", "U.S.A")} and $\kappa_{j}=\tau(q_{j})=$("Tom Cruise", "country", "U.S.A"), ("U.S.A", "continent", "North America")}. In this case, $g$ is called a superficial unlearning algorithm since $\kappa_{i}\cap\kappa_{j}\neq\text{Ø}$ and $\mathcal{M}_{\theta^{\prime}}(q_{j},a_{j})=1$ is true; thus, the equation 2 is satisfied.
@@
Based on the definition of superficial unlearning, we construct three new types of datasets: paraphrased, multi-hop, and same-answer sets to investigate the phenomenon of superficial unlearning. The paraphrased set $\mathcal{C}_{p}^{i}$, multi-hop set $\mathcal{C}_{m}^{i}$, and same-answer set $\mathcal{C}_{s}^{i}$ is matched with each question-answer pair $(q_{i},a_{i})\in\mathcal{C}$. The paraphrased set includes the same context questions with varying textual forms to the matched target question; thus, we should unlearn $\mathcal{C}_{p}^{i}$ if a matched question-answer pair $(q_{i},a_{i})$ is included in the forget set $\mathcal{C}_{f}$. The multi-hop set includes multi-hop question-answer pairs interconnected with the target question. Therefore, we should also unlearn $\mathcal{C}_{m}^{i}$ if a mapped pair $(q_{i},a_{i})$ is included in $\mathcal{C}_{f}$. The same-answer set includes question-answer pairs where the questions are from different contexts but share the same answer as $a_{i}$; thus, we should maintain the knowledge of the same-answer set, although a matched pair $(q_{i},a_{i})$ is included in $\mathcal{C}_{f}$.
@@ 3.2 Data Collection and Construction
Data Source. We construct FaithUn using Wikidata (Vrandečić and Krötzsch, 2014), a knowledge base including knowledge triples $(s,r,o)$ matched with millions of entities. We first select 200 of the most famous people as the entity set $\mathcal{E}$ from The Most Famous People Rank  [^1] https://today.yougov.com, and manually select 19 common relations as the relation set $\mathcal{R}$. The selected relations are shown in Appendix A.3.1.
The Base QA dataset. We retrieve all the triples $(s,r,o)$ from Wikidata, where $s\in\mathcal{E}$ and $r\in\mathcal{R}$. Based on these triples, we use GPT-4o mini [^2] https://openai.com/ to generate natural language form questions using a prompt template shown in Figure 6. We use an object (i.e., $o$) of each triple as the answer for each generated question. The constructed Base QA dataset $\mathcal{C}$ is split into three types of datasets: forget set $\mathcal{C}_{f}$, retain set $\mathcal{C}_{r}$, and test set $\mathcal{C}_{t}$.
Evaluation of Unlearning Generalization. We also generate the Paraphrased QA dataset $\mathcal{C}_{p}$ to evaluate the generalization of an unlearning method. Each question-answer pair $(q,a)\in\mathcal{C}$ is matched with three paraphrased questions. The Paraphrased QA dataset is generated during the Base QA dataset construction process by making GPT-4o mini generate four different questions for each triple. We use the first question as a sample of the Base QA dataset and the others for the Paraphrased QA dataset. We have strictly checked whether there are the same texts in the generated four texts by examining the lexical overlap between texts.
Evaluation of Multi-hop Knowledge Unlearning. We construct the Multi-hop QA dataset $\mathcal{C}_{m}$ to investigate superficial unlearning. Each question-answer pair $(q,a)\in\mathcal{C}$ is matched with multi-hop questions. After constructing the triples of the Base QA dataset, we additionally retrieve a set of chain-of-triples $((s_{1},r_{1},o_{1}),(s_{2},r_{2},o_{2}))$ from Wikidata, where $s_{1}\in\mathcal{E}$ and $r_{1},r_{2}\in\mathcal{R}$ and $o_{1}=s_{2}$. For each chain-of-triples, we also generate natural language questions using GPT-4o mini with the prompt template shown in Figure 7. We strictly validate that $o_{1}$ and $o_{2}$ are not included in the questions using instructions.
Evaluation of Shortcut Unlearning. We further build the Same-answer QA dataset $\mathcal{C}_{s}$. Each question-answer pair $(q,a)\in\mathcal{C}$ is also matched with the same-answer but different-context questions. After constructing the triples of the Base QA dataset, we also retrieve other triples $(s^{\prime},r^{\prime},o)$ that share the same object (i.e., $o$) with each triple from the Base QA dataset, where $s^{\prime}\notin\mathcal{E}$. We also generate natural language form questions using GPT-4o mini with the same prompt template used in constructing the Base QA dataset.

### v2

@@ Abstract
Various studies have attempted to remove sensitive or private knowledge from a language model to prevent its unauthorized exposure. However, prior studies have overlooked the inherent complexity and interconnectedness of knowledge, which requires careful examination. To resolve this problem, we first define a new concept called superficial unlearning, which refers to the phenomenon where an unlearning method either fails to erase the interconnected knowledge it should remove or unintentionally erases irrelevant knowledge. Based on the definition, we introduce a novel benchmark, FaithUn, to analyze and evaluate the faithfulness of unlearning in real-world knowledge QA settings. Furthermore, we propose a novel unlearning method, KLUE, which updates only knowledge-related neurons to achieve faithful unlearning. KLUE leverages a regularized explainability method to localize contextual knowledge neurons, updating only these neurons using carefully selected unforgotten samples. Experimental results demonstrate that existing unlearning methods fail to ensure faithful unlearning, while our method shows significant effectiveness in real-world QA unlearning.
@@
Nakyeong Yang1, Minsung Kim1, Seunghyun Yoon2, Joongbo Shin3, Kyomin Jung1 1Seoul National University, 2Adobe Research, 3LG AI Research {yny0506, kms0805, kjung}@snu.ac.kr syoon@adobe.com, jb.shin@lgresearch.ai
@@ 1 Introduction
Large language models (LLMs) are trained on a vast corpus of text, enabling them to achieve outstanding performance across various tasks. However, LLMs may present privacy risks, as sensitive or private information may be inadvertently included in the text corpus used for training. Therefore, prior studies have examined unlearning undesirable knowledge in LLMs (Shi et al., 2024; Li et al., 2024; Maini et al., 2024; Jin et al., 2024; Lynch et al., 2024; Wu et al., 2024).
@@
Figure 1: Faithful Forgetting in LLMs. FaithUn proposes three datasets to evaluate unlearning methods (i.e., Paraphrased, Multi-hop, and Same-answer datasets). Each target knowledge to be unlearned is mapped with questions from these three datasets for evaluation.
However, they are limited in that they have overlooked the complex and interconnected nature of knowledge, which necessitates a careful investigation of its internal dependencies. Figure 1 presents an example of faithful unlearning. Unlearning methods should also remove knowledge that is interconnected with the target questions to be unlearned—such as that found in paraphrased and multi-hop questions. Conversely, unlearning methods should retain knowledge that may appear relevant but is not directly connected to the target, such as questions that merely share the same answer. The unlearning process substantially relies on less data in training, as its goal is to remove only specific knowledge. Therefore, unlearned models tend to collapse into trivial solutions, unlike general training that utilizes large-scale data and enables broad generalization and multi-hop reasoning.
@@
   Knowledge Exists in LLMs
   Generalization Test
   Multi-hop Unlearning Test
   Shortcut Unlearning Test
@@
Table 1: Dataset Comparison. FaithUn aims to examine three challenges: generalization, multi-hop knowledge unlearning, and shortcut unlearning to investigate superficial unlearning. FaithUn can be used flexibly to evaluate unlearning methods since it targets pre-existing knowledge of famous figures within LLMs.
To address these problems, we first define superficial unlearning, which refers to the phenomenon where an unlearning method either fails to erase the interconnected knowledge it should remove or unintentionally erases irrelevant knowledge. Based on the definition, we introduce FaithUn (Faithful Unlearning Evaluation Benchmark), a new benchmark to investigate superficial unlearning. We construct three datasets—paraphrased, multi-hop, and same-answer—each addressing a key challenge: generalization, multi-hop knowledge unlearning, and shortcut unlearning, respectively. We demonstrate that existing unlearning methods do not ensure faithful unlearning, which raises new research questions for knowledge unlearning.
Furthermore, we propose a robust method, KLUE (Knowledge-Localized UnlEarning) to achieve faithful unlearning by precisely identifying and updating neurons related to the target knowledge. Specifically, we use the attribution method (Yang et al., 2023) to determine which neurons should be updated by quantifying how much each neuron contributes to predicting the answer to a given question. However, the quantified score may include superficial knowledge that simply affects the target output’s probability without considering contextual meaning. Therefore, we propose a novel knowledge regularization method that accurately quantifies each neuron’s knowledge score, mitigating the trivial contribution of neurons. After identifying knowledge neurons, our method selectively unlearns the target knowledge while preserving other knowledge by updating only knowledge-related neurons with selected unforgotten samples. Our experiments reveal that existing methods fail to ensure faithful unlearning. However, KLUE significantly outperforms the baselines in the FaithUn setting, demonstrating that knowledge-localized unlearning effectively achieves faithful unlearning. In summary, this work makes the following contributions:
  • We define superficial unlearning and introduce FaithUn, a new benchmark for evaluating whether unlearning methods can faithfully handle the interconnectedness of world knowledge.
  • We reveal that existing methods fail to achieve faithful unlearning by showing a trivial solution, highlighting the need for further research.
  • We propose KLUE, a knowledge-localized unlearning method that regularizes neuron attribution to identify and selectively update context-relevant neurons, achieving superior performance on FaithUn.
2 Large Language Models Unlearning
Machine unlearning has been used as a solution to address privacy and copyright issues in the text generation process of LLMs. Notable examples include gradient ascent-based methods (Jang et al., 2023; Yao et al., 2023; Barbulescu and Triantafillou, 2024), preference optimization approaches (Rafailov et al., 2024; Zhang et al., 2024; Jin et al., 2024), and representation learning techniques (Li et al., 2024; Yao et al., 2024). However, the effectiveness of these methods has not been clearly demonstrated, prompting the introduction of new benchmarks in the unlearning field to assess them. WHP, MUSE, and KnowUndo (Eldan and Russinovich, 2023; Shi et al., 2024; Tian et al., 2024) have aimed to unlearn the knowledge of copyrighted texts (e.g., News and Book). WMDP (Li et al., 2024) has introduced a benchmark for hazardous knowledge in professional domains (e.g., biosecurity). TOFU (Maini et al., 2024) has created synthetic profiles and removed their associated knowledge from LLMs. RWKU (Jin et al., 2024) has examined knowledge about real-world entities and evaluates memorization across various textual forms (e.g., cloze tests and QA) to assess generalization. While these studies have made valuable contributions, they fail to address the interconnected nature of knowledge. Even RWKU, despite its progress in surface-level generalization, fails to capture the deeper relational dependencies among pieces of knowledge. As a result, prior works have overlooked two critical issues—knowledge interconnections and shortcut unlearning—which are essential due to the inherently limited data available for unlearning. We summarize the differentiations of our benchmark compared to others in Table 1. Furthermore, we provide the detailed dataset comparisons in Appendix A.1.
3 The FaithUn Benchmark
3.1 Problem Definition
The FaithUn task evaluates unlearning algorithms under real-world knowledge QA settings. Formally, given a language model $P_{\theta}(y|x)=\prod_{t=1}^{T}P_{\theta}(y_{t}|x,y_{1},...,y_{t-1})$ with parameters $\theta$, an unlearning algorithm $f$ updates $\theta$ to $\theta^{\prime}$, erasing the target knowledge from $P_{\theta}$. FaithUn includes various question-answer pairs $(q,a)\in\mathcal{C}$, where $\mathcal{C}$ is a question-answer pair set. Our task provides forget set $\mathcal{C}_{f}\subset\mathcal{C}$, which contains target question-answer pairs to be forgotten. FaithUn also provides retain set $\mathcal{C}_{r}\subset\mathcal{C}\backslash\mathcal{C}_{f}$ and test set $\mathcal{C}_{t}\subset\mathcal{C}\backslash(\mathcal{C}_{f}\cup\mathcal{C}_{r})$. $\mathcal{C}_{r}$ is used in the unlearning process as training samples to maintain the original knowledge of $P_{\theta}$, and $\mathcal{C}_{t}$ is used as unseen data to evaluate an unlearned model $P_{\theta^{\prime}}$ to determine whether the unlearned model maintains the original knowledge. Furthermore, FaithUn provides other new types of test sets (i.e., paraphrased, multi-hop, and same-answer sets) to assess the faithfulness of unlearning methods. Before introducing the other datasets, we first define key aspects of our benchmark.
@@ 3.1 Problem Definition
A world knowledge graph $\mathcal{K}$ is a directed multi-graph where nodes are entities and edges are labeled with relations, i.e., elements of two sets $\mathcal{E}$ and $\mathcal{R}$, respectively. We define $\mathcal{K}$ as a collection of triples $(s,r,o)\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$, where $s$, $r$, $o$ denote the subject, relation, and object, respectively (Ruffinelli et al., 2020). We assume that a world knowledge question is mapped to triples of $\mathcal{K}$; thus, we also define a knowledge mapping function, $\tau:\mathcal{Q}\rightarrow\mathcal{P}(\mathcal{K})$, where $\mathcal{Q}$ is a set of questions and $\mathcal{P}(\mathcal{K})$ represents the power set of $\mathcal{K}$. For example, the knowledge of a multi-hop question, $q_{i}$ = "Which continent is Tom Cruise’s country in?", can be denoted as a set of triples like $\kappa_{i}=\tau(q_{i})$ = {("Tom Cruise", "country", "U.S.A"), ("U.S.A", "continent", "North America")}.
@@
where $\iota$ is an input prompt template for the language model $P_{\theta}$, and $\mathcal{Q}$ and $\mathcal{A}$ are question and answer sets. Therefore, $\mathcal{M}_{\theta}(q,a)=1$ indicates that the model retains the knowledge of $(q,a)$, while $\mathcal{M}_{\theta}(q,a)=0$ signifies that it does not.
@@
where $\kappa_{f}=\bigcup_{(q,a)\in\mathcal{C}_{f}}{\hskip-1.42271pt\tau(q)}$ and $\kappa_{j}=\tau(q_{j})$.
For example, suppose that an unlearning algorithm $g$ unlearns the question $q_{i}$ = "Which country is Tom Cruise from?", but it does not unlearn the multi-hop question $q_{j}$ = "Which continent is Tom Cruise’s country in?". Then, the knowledge of two questions can be denoted as a set of knowledge triples like $\kappa_{i}=\tau(q_{i})=$ {("Tom Cruise", "country", "U.S.A")} and $\kappa_{j}=\tau(q_{j})=$ {("Tom Cruise", "country", "U.S.A"), ("U.S.A", "continent", "North America")}. In this case, $g$ is called a superficial unlearning algorithm since $\kappa_{i}\cap\kappa_{j}\neq\text{Ø}$ and $\mathcal{M}_{\theta^{\prime}}(q_{j},a_{j})=1$ is true (11st condition).
In another case, suppose $g$ unlearns only $\kappa_{i}$ = {("Tom Cruise", "country", "U.S.A")} but mistakenly also remove $\kappa_{j}$ = {("Andy Warhol", "country", "U.S.A")}. This satisfies superficial unlearning since $\kappa_{i}\cap\kappa_{j}=\text{Ø}$ and $\mathcal{M}_{\theta^{\prime}}(q_{j},a_{j})=0$ (22nd condition).
@@
Based on the definition of superficial unlearning, we construct three new types of datasets: paraphrased, multi-hop, and same-answer sets to investigate the phenomenon of superficial unlearning. The paraphrased set $\mathcal{C}_{p}^{i}$, multi-hop set $\mathcal{C}_{m}^{i}$, and same-answer set $\mathcal{C}_{s}^{i}$ is matched with each question-answer pair $(q_{i},a_{i})\in\mathcal{C}$. The paraphrased set includes the same context questions with varying textual forms to the matched target question; thus, we should unlearn them if a matched question-answer pair $(q_{i},a_{i})$ is included in the forget set. The multi-hop set includes multi-hop question-answer pairs interconnected with the target question. Therefore, we should also unlearn them if a mapped pair $(q_{i},a_{i})$ is included in the forget set. The same-answer set includes question-answer pairs where the questions are from different contexts but share the same answer as $a_{i}$; thus, we should maintain the knowledge of the same-answer set, although a matched pair $(q_{i},a_{i})$ is included in the forget set.
@@ 3.2 Data Collection and Construction
Data Source. We construct FaithUn using Wikidata (Vrandečić et al., 2014), a knowledge base including knowledge triples $(s,r,o)$ matched with millions of entities. We first select 200 of the most famous people as the entity set $\mathcal{E}$ from The Most Famous People Rank  [^1] https://today.yougov.com, and manually select 19 common relations as the relation set $\mathcal{R}$. The selected relations are shown in Appendix A.2.1.
The Base QA dataset. We retrieve all the triples $(s,r,o)$ from Wikidata, where $s\in\mathcal{E}$ and $r\in\mathcal{R}$. Based on these triples, we use GPT-4o mini to generate natural language form questions using a prompt template shown in Figure 6. Note that converting triples into natural language questions is a simple task, and most LLMs are capable of performing it. We use an object (i.e., $o$) of each triple as the answer for each generated question. The constructed Base QA dataset is split into three types of datasets: forget set, retain set, and test set.
Assessing Unlearning Generalization. We also generate the Paraphrased QA dataset to evaluate the generalization of an unlearning method. Each question-answer pair $(q,a)\in\mathcal{C}$ is matched with three paraphrased questions. The Paraphrased QA dataset is generated during the Base QA dataset construction process by making GPT-4o mini generate four different questions for each triple. We use the first question as a sample of the Base QA dataset and the others for the Paraphrased QA dataset. We have strictly checked whether there are the same texts in the four generated texts by examining the lexical overlap between texts.
Assessing Multi-hop Unlearning. We construct the Multi-hop QA dataset to investigate superficial unlearning. Each question-answer pair $(q,a)\in\mathcal{C}$ is matched with multi-hop questions. After constructing the triples of the Base QA dataset, we additionally retrieve a set of chain-of-triples $((s_{1},r_{1},o_{1}),(s_{2},r_{2},o_{2}))$ from Wikidata, where $s_{1}\in\mathcal{E}$ and $r_{1},r_{2}\in\mathcal{R}$ and $o_{1}=s_{2}$. For each chain-of-triples, we also generate questions using GPT-4o mini with the prompt template shown in Figure 7. We ensure that $o_{1}$ and $o_{2}$ are not included in the questions with an instruction, and validate this with the lexical overlaps.
Assessing Shortcut Unlearning. We further build the Same-answer QA dataset. Each question-answer pair $(q,a)\in\mathcal{C}$ is also matched with the same-answer but different-context questions. After constructing the triples of the Base QA dataset, we also retrieve other triples $(s^{\prime},r^{\prime},o)$ that share the same object (i.e., $o$) with each triple from the Base QA dataset, where $s^{\prime}\notin\mathcal{E}$. We also generate questions using GPT-4o mini with the same prompt template used in constructing the Base QA dataset.